Find the deflection of the fixed end vibrating string of unit length corresponding to zero initial deflection and u(x) given below as the initial velocity.

x , 0 ≤ x < 1/2
u(x) =
1-x , 1/2 ≤ x <1
1
Expert's answer
2019-09-02T10:37:18-0400

Let "w(x,t)" be the function describing deflection of the string. The function satisfies wave equation

"w_{tt}=a^2w_{xx}"

and from the condition of the question we have

"w(0,t)=w(1,t)=0"

as the ends are fixed and initial conditions are


"w(x,0)=0"


"w_t(x,0)=u(x)"

In order to solve the PDE we’ll use separation of variables, so "w(x,t)=X(x)T(t)". If we put this in the equation, we’ll obtain "\frac{T''}{a^2T}=\frac{X''}{X}=-\lambda". So we have Sturm-Liouville problem "\begin{cases}
X''+\lambda X=0\\
X(0)=X(1)=0
\end{cases}". The problem has infinitely many non-trivial solutions "X_n=\sin(\pi n x)" and eigenvalues "\lambda_n=\pi^2n^2". Then "T_n''+ \lambda_n a^2 T_n=0" hence "T_n=C_{1n}\cos(\pi n at)+C_{2n}\sin(\pi n a t)" where "C_{1n}, C_{2n}" are arbitrary constants. So

"w(x,t)=\sum_{n=1}^{\infty}T_n(t)X_n(x)=\sum_{n=1}^{\infty}(C_{1n}\cos(\pi n at)+C_{2n}\sin(\pi n a t))\sin(\pi n x)"


In order to find "C_{1n}, C_{2n}" we’ll use initial conditions. "w(x,0)=\sum_{n=1}^{\infty}C_{1n}\sin(\pi n x)=0", so all "C_{1n}=0"

"w_t(x,0)=\sum_{n=1}^{\infty}C_{2n}\pi n a\sin(\pi n x)=u(x)"


let’s expand "u(x)" into Fourier sine series on interval "0<x<1".


"u(x)=\sum_{n=1}^{\infty}b_n \sin(\pi n x)"


"b_n=2 \int_0^{1} u(x) \sin(\pi n x)\, dx=\frac{4\sin(\frac{\pi n}{2})}{\pi^2 n^2}"

So


"u(x)=\sum_{n=1}^{\infty}\frac{4\sin(\frac{\pi n}{2})}{\pi^2 n^2} \sin(\pi n x)"

and equating coefficients in the sums we obtain "C_{2n}=\frac{4\sin(\frac{\pi n}{2})}{\pi^3n^3a}". So


"w(x,t)=\sum_{n=1}^{\infty}\frac{4\sin(\frac{\pi n}{2})}{\pi^3n^3a}\sin(\pi n a t)\sin(\pi n x)"

or it can be rewritten as 


"w(x,t)=\sum_{n=1}^{\infty}(-1)^{n+1}\frac{4}{\pi^3(2n-1)^3a}\sin(\pi (2n-1) a t)\sin(\pi (2n-1) x)"